OpenStudy (jamanorthi):

write a formula for finding a midpoint of a segment when one of the endpoints is the origin

OpenStudy (anonymous):

if other end is (x,y), then mid point=(x/2,y/2)

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ \square }}\quad ,&{\color{blue}{ \square }})\quad &({\color{red}{ \square }}\quad ,&{\color{blue}{ \square }}) \end{array}\qquad \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\\ \qquad \quad \uparrow \\ \textit{origin, thus replace with 0,0}\)

OpenStudy (jamanorthi):

so the formula is that jdoe0001

OpenStudy (jamanorthi):

ok thanks guys